# Do Unlearning Methods Remove Information from Language Model Weights?

### v1

@@ 1 Introduction
The evaluations of unlearning methods are mostly output-based, which fails to determine if the knowledge is removed from the model weights. Lynch et al. (2024) showed that even after applying the unlearning method suggested by Eldan & Russinovich (2023), information could be recovered from the model using multiple methods, including simply changing the format of questions. Even when applying RMU (Li et al., 2024b) (a state-of-the-art unlearning technique that targets removing harmful knowledge), harmful information can still be recovered using jailbreaks (Li et al., 2024a). To develop reliable unlearning methods, we need to develop robust evaluations to guide the research process.
@@ 2 Related Work
Previous work has measured the success of unlearning using performance on a task related to the unlearned information, or output similarity to that of a model that was not trained on the information to be unlearned (Nguyen et al., 2022; Lynch et al., 2024; Liu et al., 2024a), but these approaches measure the propensity of the LLM to use the unlearned knowledge, failing to capture hidden knowledge. Liu et al. (2024a) suggests two metrics to assess unlearning effectiveness: evaluating unlearning on harder cases (e.g., jailbreaks, queries in different languages) and Membership Inference Attacks (MIA) (Shokri et al., 2016). Since it is not possible to try all jailbreaks, evaluating jailbreak robustness is difficult: even if some attacks fail, others may succeed. For example, Li et al. (2024a) demonstrates that RMU (Li et al., 2024b) could be jailbroken using hand-crafted attacks, despite its high robustness against many automated attacks. MIA do not measure the absence of knowledge about a particular fact, but the likelihood that a particular datapoint is absent from the training corpus, which is not the relevant metric for the purpose of preventing LLM misuse.
@@
Past work has introduced white-box unlearning evaluations like linear probes and relearning time. Linear probes may recover the information present in the activations (Lynch et al., 2024), but are not powerful enough to detect information present in the weights. For example, probes on top of RMU models fail to get high accuracy, but RMU models can still be jailbroken. Relearning time is a promising approach to evaluate unlearning that was used by Tarun et al. (2023) and Lynch et al. (2024). This approach measures the number of fine-tuning steps needed to recover accuracy on the unlearned knowledge, where the fine-tuning data is drawn from the dataset we perform unlearning on. If accuracy is recovered after a small number of fine-tuning steps, this is evidence that the unlearning technique is ineffective at removing knowledge. The main problem with this approach is that it produces a metric that is only meaningful when compared to the learning time for a model that was never trained on the knowledge in the first place, which is hard to obtain given the cost of LLM pretraining. This method also suffers against methods that specifically target making fine-tuning slower or more difficult (Rosati et al., 2024; Henderson et al., 2023).
@@ Appendix B RTT Loss
Appendix C Testing Against Robust Unlearning Methods

### v2

@@ 1 Introduction
The evaluations of unlearning methods are mostly output-based, which fails to determine if the knowledge is removed from the model weights. Lynch et al. (2024b) showed that even after applying the unlearning method suggested by Eldan & Russinovich (2023), information could be recovered from the model using multiple methods, including simply changing the format of questions. Even when applying RMU (Li et al., 2024b) (a state-of-the-art unlearning technique that targets removing harmful knowledge), harmful information can still be recovered using jailbreaks (Li et al., 2024a). To develop reliable unlearning methods, we need to develop robust evaluations to guide the research process.
@@ 2 Related Work
Previous work has measured the success of unlearning using performance on a task related to the unlearned information, or output similarity to that of a model that was not trained on the information to be unlearned (Nguyen et al., 2022; Lynch et al., 2024b; Liu et al., 2024a), but these approaches measure the propensity of the LLM to use the unlearned knowledge, failing to capture hidden knowledge. Liu et al. (2024a) suggests two metrics to assess unlearning effectiveness: evaluating unlearning on harder cases (e.g., jailbreaks, queries in different languages) and Membership Inference Attacks (MIA) (Shokri et al., 2016). Since it is not possible to try all jailbreaks, evaluating jailbreak robustness is difficult: even if some attacks fail, others may succeed. For example, Li et al. (2024a) demonstrates that RMU (Li et al., 2024b) could be jailbroken using hand-crafted attacks, despite its high robustness against many automated attacks. MIA do not measure the absence of knowledge about a particular fact, but the likelihood that a particular datapoint is absent from the training corpus, which is not the relevant metric for the purpose of preventing LLM misuse.
@@
Past work has introduced white-box unlearning evaluations like linear probes and relearning time. Linear probes may recover the information present in the activations (Lynch et al., 2024b), but are not powerful enough to detect information present in the weights. For example, probes on top of RMU models fail to get high accuracy, but RMU models can still be jailbroken. Relearning time is a promising approach to evaluate unlearning that was used by Golatkar et al. (2020a), Golatkar et al. (2020b), Tarun et al. (2023), and Lynch et al. (2024b). This approach measures the number of fine-tuning steps needed to recover accuracy on the unlearned knowledge, where the fine-tuning data is drawn from the dataset we perform unlearning on. If accuracy is recovered after a small number of fine-tuning steps, this is evidence that the unlearning technique is ineffective at removing knowledge. The main problem with this approach is that it produces a metric that is only meaningful when compared to the learning time for a model that was never trained on the knowledge in the first place, which is hard to obtain given the cost of LLM pretraining. This method also suffers against methods that specifically target making fine-tuning slower or more difficult (Henderson et al., 2023; Rosati et al., 2024; Tamirisa et al., 2024).
Weaknesses of current unlearning techniques
Previous work has shown evidence about current unlearning techniques being weak against attacks that would fail if the information was removed (Lynch et al., 2024a; Łucki et al., 2024; Hong et al., 2024). Our results further confirm the findings of this previous work and provides a standard method that can be applied to any unlearned model (a model to which unlearning was applied). Further, negative results using our method provide strong evidence of unsuccessful knowledge removal relative to previous prompting-based methods.
@@ 8 Conclusion
Acknowledgments
The authors would like to thank Buck Shlegeris, Lawrence Chan, Robert Kirk, and Xander Davies for help and feedback on this paper, the ML Alignment Theory Scholars (MATS) program for their support, and a lot of other people for helpful discussions of these ideas.
@@ Appendix B RTT Loss
Appendix C Fine-tuning Robust Methods
@@
Other techniques include Tampering Attack Resistance (TAR) introduced by Tamirisa et al. (2024), but this technique is vulnerable against parameter-efficient fine-tuning (PEFT) as demenstorated in the work.
Overall, because fine-tuning robustness techniques can be bypassed when using an extensive hyperparameter search, we think using RTT with an extensive hyperparameter search would still expose knowledge that wasn’t removed.